elimination problem: two angles are complementary. the sum of the measure of the first angle and half the second angle is 64. find the measure of the smaller angle
Answer by checkley77(12844)

X+Y/2=64
-(X+Y=90) SUBTRACT
-------------------
Y/2-Y=64-90
-Y/2=-26 CROSS MULTIPLY.
-Y=-26*2
-Y=-52
Y=52 DEGREES FOR THE LARGER ANGLE.
X+52/2=64
X+26=64
X=64-26
X=38 DEGREES FOR THE SMALLER ANGLE (X).
PROOF:
38+52=90
90=90